OpenStudy (anonymous):

A car is initially moving with a velocity of 3 m/s. then it accelerates uniformly at the rate of 0.6 m/s^2 for 5 seconds. What is the distance traveled by the car in 5 seconds?

OpenStudy (anonymous):

\[s = V_o*t + \frac{1}{2}*a*t^2 \]

OpenStudy (anonymous):

ahhh I dont understand this stuff.

OpenStudy (anonymous):

\[s = 3 \times 5 + \frac{1}{2}\times 5^2 \times .6\] \[s = 15 + \frac{1}{2}\times 25\times \frac{6}{10}\]

OpenStudy (anonymous):

this will give me the distance?

OpenStudy (anonymous):

I don't understand it either but I keep practicing, you must practice and keep trying; you will understand all this soon. Yes, It will

OpenStudy (anonymous):

i am seriously behind. like my labs, I am like 6 weeks behind >.< due to not understanding! would you mind helping me with my quiz?

OpenStudy (anonymous):

I will try my best! I am so behind too only 5 months left and I have to complete Mechanics, Thermodynamics, Optics, Electrostats and Modern Physics, Life hits you hard sometime

OpenStudy (anonymous):

22.5

OpenStudy (anonymous):

holy crap kidddo! :) you very sciencey...not a word, i know.. but WOW! I would really appreciate teh help :)

OpenStudy (anonymous):

bharatindians! That was right!!! Thank you :))

OpenStudy (anonymous):

if u have problems simple concepts then try this dear.. that will really helps u..

OpenStudy (anonymous):

 so thevertical is position(x) and horizontal is time (sec) its asking if it represents uniform velocity, unirform accerleration, the objesct is at rest , or its accelerationg and decelerating..I have no idea?

OpenStudy (anonymous):

HELP! PLEASE!?

OpenStudy (anonymous):

http://www.khanacademy.org/ that's really help u in ur studies

OpenStudy (anonymous):

accelerating according to me

OpenStudy (anonymous):

thats what i thought, nbut idk what any of the uniform ones mean or what not :/

OpenStudy (anonymous):

oh wait it must be uniform acceleration

OpenStudy (anonymous):

:) imma just post the quiz stuff here kay?

OpenStudy (anonymous):

the graph is position time I assumed it as velocity time silly me sorry it must be uniform acceleration yeah it's fine with me

OpenStudy (anonymous):

What does the area under the velocity-time graph represent? (Points : 1) acceleration displacement speed velocity

OpenStudy (anonymous):

displacement, pretty sure this time

OpenStudy (anonymous):

i said volcity :/

OpenStudy (anonymous):

oh my bad =P

OpenStudy (anonymous):

An object is thrown with an initial velocity of 3 m/s. What is its acceleration at its maximum height? (Points : 2) zero 3 m/s2 9.8 m/s2 none of the above I saidzero :/

OpenStudy (anonymous):

umm I think it's 9.8

OpenStudy (anonymous):

Gravity Works all the time